A truck enters a highway driving 60mph. a car enters the highway at the same spot, 8 minutes later and drives 71 mph in the same direction. From the time the car enters the highway, how long will it take the car to pass the truck?
Answer by checkley77(12844)

D=RT
D=60T FOR THE TRUCK.
D=71(T-8/60)
60T=71(T-4/30)
60T=71T-284/30
60T-71T=-9.467
-11T=-9.467
T=-9.467/-11
T=.86 HOURS OR 51.64 MINUTES TO PASS THE TRUCK.
PROOF:
60*.86=71(.86-8/60
51.6=71(.86-.1333)
51.6=71*.627
51.6=51.6